A rescue pilot drops a survival kit while her plane is flying at an altitude of 1200.0 meters with a forward velocity of 97.4 m/s. How far in advance of the drop zone should she release the package?


1
Expert's answer
2020-12-17T07:37:40-0500

Time of falling "t" can be found from the kinematic equation (if there is no vertical initial velocity component):


"h = \dfrac{gt^2}{2}"

where "h = 1200m" is the altitude, and "g = 9.81m\/s^2" is the gravitational acceleration. Thus, obtain:


"t = \sqrt{\dfrac{2h}{g}} = \sqrt{\dfrac{2\cdot 1200}{9.81}} \approx 15.6s"

In this time the survival kit covers the following horizontal distance:


"d = v_0t = 97.4m/s\cdot 15.6s \approx 1523m"

Here "v_0 = 97.4m\/s" is the initial horizontal velocity of the kit. Thus, it should be release 1523 m in advance.


Answer. 1523 m.